# Imbalanced Data Clustering using Equilibrium K-Means

### v2

@@ VI Deep Clustering
Clustering algorithms using Euclidean distance for defining similarity face a challenge when clustering high-dimensional data, such as images, due to the small distance differences between various point pairs in high-dimensional space [41]. Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data to low-dimensional representation. Many deep clustering methods combine DNNs with HKM, e.g., [15, 16, 17]. However, this design is less favorable to imbalanced data [16]. In this section, we show the promise of EKM in the deep clustering of imbalanced data by replacing HKM with EKM in a popular deep clustering framework.
@@ VI-A Optimization Procedure
where $m_{k}^{n}$ is the number of samples assigned to the $k$-th cluster before the incoming data $\mathbf{x}_{n}$, controlling the learning rate of the $k$-th centroid. Overall, the optimization procedure of DCN alternates between updating networks parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (45) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (46) and (47).
@@ VI-B Clustering Performance on MNIST
To implement DCN, we refer to the code one of its authors provided, available at https://github.com/boyangumn/DCN-New. We use the default neural network structure and hyperparameters. In particular, the dimension of the low-dimensional representation is set to ten, and the parameter $\beta$ is set to one. The smoothing parameter $\alpha$ is tuned on the representation obtained by the initialized DCN networks according to the strategy introduced in Section IV-C. We evaluate clustering performance using three widely accepted metrics: including normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC). NMI and ACC range from 0 to 1, with zero representing the worst and one representing the best. ARI ranges from -1 to 1, with minus one representing the worst and one representing the best.
We first evaluate the algorithm’s performance on the full MNIST dataset [42], which contains 70,000 gray images of handwritten digits from 0 to 9. Each image has $28\times 28=784$ pixels. We set the smoothing parameter of EKM to $\alpha=5\mathrm{e}-3$. The clustering results are presented in TABLE I. We compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE). SAE is a specific version of DCN that only minimizes the reconstruction error. Thus, the learned representation by SAE does not have a clustering-friendly structure. The results show that DCN outperforms SAE, highlighting the importance of a clustering-friendly structure. We can also see that DCN+EKM performs similarly to DCN+HKM. We omit the results of DCN+FKM and DCN+MEFC due to space limitations. Their performance is not better than DCN+HKM.
MNIST contains 60,000 training images and 10,000 testing images. To evaluate the performance of imbalanced data clustering, we generate an imbalanced MNIST dataset by removing the training images of digits 1 to 9. This imbalanced dataset contains 15,923 images, of which approximately 6,900 are digit 0, while the remaining digits (1 to 9) each have about 1,000 images. We set the smoothing parameter of EKM to $\alpha=3.8\mathrm{e}-3$. The results are summarized in TABLE II. In conclusion, we find that an EKM-friendly representation structure is crucial for deep clustering of imbalanced data. We map the ten-dimensional representation obtained by DCN to a two-dimensional space by t-SNE [43] for visualization. The results are displayed in Fig. 12. The visualization result implies that the deep representation obtained by DCN+EKM is more discriminative than that obtained by DCN+HKM. We observe that EKM successfully identifies the large class (number 0) from other small classes without referring to the true labels, while HKM incorrectly divides the large class into four clusters.
TABLE I: Evaluation on full MNIST
@@
TABLE II: Evaluation on imbalanced MNIST
@@
Figure 12: Visualization of DCN representations by t-SNE. (a) Representations obtained by DCN+HKM, colored by true labels. (b) Representations obtained by DCN+HKM, colored by labels generated by HKM. (c) Representations obtained by DCN+EKM, colored by true labels. (d) Representations obtained by DCN+EKM, colored by labels generated by EKM.
@@ VII Conclusion
This paper presents equilibrium K-means (EKM), a novel clustering algorithm effective for both balanced and imbalanced data. EKM is simple, interpretable, scalable to large datasets, and capable of creating unequal-sized clusters. Experimental results on real datasets from various domains show that EKM outperforms HKM, FKM, and other popular centroid-based algorithms on datasets with imbalanced data and performs comparably on datasets with balanced data. In the same deep clustering framework, compared with HKM and FKM, the clustering accuracy of imbalanced data using EKM has been noticeably improved. Furthermore, we reformulate HKM, FKM, and EKM in a general form of gradient descent. We encourage readers to study more properties of this general form to develop new K-means algorithms with non-heuristic memberships to address different issues.
@@ Appendix A Proof of Theorem 1
For concise and tidy, we denote the partial derivative $\frac{\partial h}{\partial d_{k}}(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$ as $\frac{\partial h}{\partial d_{k}}|_{n,{\tau}}$. Since $h$ is a concave function at $[0,\,+\infty)^{K}$, we have
$h(d_{1n},\cdots,d_{Kn})$ (49)
$\leq h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)}),$
@@
$J(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})\\ \leq\sum_{n=1}^{N}\bigg{[}h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)})\bigg{]}\\ =J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})+\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\bigg{(}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot\big{(}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}\\ -\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\big{)}\bigg{)}.$ (50)
@@
Supplemental Materials: Imbalanced Data Clustering using Equilibrium K-Means
supplement
We evaluate the clustering performance of the proposed equilibrium K-means (EKM) on ten real datasets. Four datasets (Iris, Imbalanced Iris, Wisconsin Diagnostic Breast Cancer (WDBC), and Imbalanced WDBC) are introduced in detail in the main text. Herein we supplement the results for the remaining six datasets. They are from the UCI repository (available at https://archive.ics.uci.edu/datasets), which are: Wine, Ecoli, Image Segmentation (IS), Htru2, Rice, and Dry Bean. A detailed description of all ten datasets are summarized in TABLE SI, ordered according to the imbalance level measured by CV0, which we will define later. The dataset with a lower order has a higher CV0 value, indicating a higher imbalance level. In the “Feature” column of TABLE SI, the format “x (y)” indicates that the dataset has y features, of which x features are used for clustering. For instance, the Iris dataset has four features, of which two are used for clustering. Partial features are used for the sake of visualizing clustering results. In the “CV0” column, we calculate the coefficient of variation (CV) of the dataset to measure the level of data imbalance. CV is defined as the ratio of the standard deviation to the mean. Mathematically, given the number of instances in each class as $N_{1},\cdots,N_{K}$, we have
$\text{CV}=\text{std}(\mathbf{N})/\text{mean}(\mathbf{N}),$
where
$\text{mean}(\mathbf{N})=\frac{\sum_{k=1}^{K}N_{k}}{K},$
and
$\text{std}(\mathbf{N})=\sqrt{\frac{\sum_{k=1}^{K}(N_{k}-\text{mean}(\mathbf{N}))^{2}}{K-1}}.$
In general, the larger the CV value, the greater the imbalance of the data.
Each dataset is normalized so that each feature has zero mean and unit variance. The number of clusters is assumed to be equal to the number of classes for all datasets. For the supplementary six datasets, the smoothness parameter $\alpha$ is chosen adaptively as follows:
$\alpha=2/\bar{d}_{0n},$
where $\bar{d}_{0n}=\frac{1}{2}\sum_{n=1}^{N}\|\mathbf{x}_{n}\|_{2}^{2}/N$, $\mathbf{x}_{n}$ is the $n$-th data point, and $N$ is the total number of data points. The value $\bar{d}_{0n}$ can be interpreted as a data variance after the normalization of zero data mean. The normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC) of each algorithm on the ten datasets are presented in TABLE SII. Bold values indicate the best results. We also include three metrics: “CV1”, ”DCV” and ”Running time” for readers’ interest. “CV1” is the CV value of the clustering result. “DCV” is the difference between CV0 and CV1, i.e., $\text{DCV}=\text{CV}0-\text{CV}1$. A positive DCV value indicates that the data distribution of the true classes is more imbalanced than that of the generated clusters, while a negative DCV value means the data distribution of the true classes is more uniform than that of the generated clusters. “Running time” is the sum of the running time of ten replicates.
We observe that the proposed EKM generally outperforms the other six algorithms on the last four datasets with CV values greater than 0.5. EKM outperforms other algorithms on two of the top six datasets with CV values less than 0.5. On the rest of the datasets, EKM still has a competitive performance. Overall, these results demonstrate that EKM is an efficient clustering algorithm for both balanced and imbalanced data.
TABLE SI: Detailed description of REAL-WORLD Datasets
Dataset Instance Feature Class CV0
Iris 150 2 (4) 3 0
IS 2310 19 (19) 7 0
Wine 178 13 (13) 3 0.1939
Rice 3810 7 (7) 2 0.2042
WDBC 569 3 (30) 3 0.3604
Dry Bean 13611 16 (16) 7 0.4951
Imbalanced Iris 120 2 (4) 2 0.9428
Ecoli 336 7 (7) 8 1.1604
Imbalanced WDBC 369 3 (30) 2 1.3222
Htru2 17898 8 (8) 2 1.4142
TABLE SII: Experimental results of different algorithms on real-world datasets
Dataset Measurement HKM FKM MEFC csiFKM siibFKM GMM EKM
Iris NMI 0.5945 0.5671 0.5194 0.5671 0.5675 0.6775 0.5457
ARI 0.5462 0.5328 0.5291 0.5328 0.5281 0.5354 0.5134
ACC 0.7867 0.7800 0.7933 0.7800 0.7800 0.7000 0.7733
CV1 0.0346 0.0200 0.1217 0.0200 0.1442 0.8902 0.0346
DCV -0.0346 -$0$0.1217 -$0$0.1442 -0.8902 -0.0346
Running time 0.0113 0.0111 0.0305 0.0142 0.1844 0.0715 0.0328
IS NMI 0.5864 0.5953 0.6222 0.5971 0.5997 0.5657 0.6463
ARI 0.4605 0.4973 0.4995 0.4962 0.3214 0.4456 0.5161
ACC 0.5455 0.6455 0.5944 0.6433 0.5861 0.5641 0.5944
CV1 0.8246 0.2450 0.6433 0.4617 1.0377 0.5037 0.6582
DCV -0.8246 -0.2450 -0.6433 -0.4617 -1.0377 -0.5037 -0.6582
Running time 0.0321 0.4714 0.2616 0.7913 49.5369 1.0463 0.7205
Wine NMI 0.8759 0.8759 0.8759 0.8610 0.3361 0.8466 0.8920
ARI 0.8975 0.8975 0.8975 0.8804 0.1355 0.8649 0.9134
ACC 0.9663 0.9663 0.9663 0.9607 0.5562 0.9551 0.9719
CV1 0.1242 0.1242 0.1242 0.1219 1.1513 0.1070 0.1403
DCV 0.0696 0.0696 0.0696 0.0720 -0.9574 0.0868 0.0535
Running time 0.0075 0.0153 0.0088 0.0126 1.6469 0.0515 0.0094
Rice NMI 0.5685 0.5688 0.5699 0.5554 0.5415 0.4891 0.5653
ARI 0.6815 0.6824 0.6833 0.6642 0.6394 0.5649 0.6772
ACC 0.9129 0.9131 0.9134 0.9076 0.9000 0.8759 0.9115
CV1 0.2442 0.2183 0.2309 0.2977 0.3623 0.0163 0.2643
DCV -0.0401 -0.0141 -0.0267 -0.0935 -0.1581 0.1878 -0.0601
Running time 0.0200 0.0438 0.0357 0.1032 2.1275 0.2303 0.0397
WBDC NMI 0.4920 0.4870 0.5005 0.4423 0.2951 0.4618 0.4906
ARI 0.5844 0.5963 0.6257 0.4402 0.2797 0.5722 0.5340
ACC 0.8840 0.8875 0.8963 0.8366 0.7750 0.8787 0.8682
CV1 0.5990 0.5393 0.4051 0.8227 0.9768 0.2461 0.7133
DCV -0.2386 -0.1790 -0.0447 -0.4623 -0.6164 0.1143 -0.3529
Running time 0.0156 0.0226 0.0171 0.0385 1.7632 0.0894 0.0160
Dry Bean NMI 0.7138 0.7055 0.7050 0.6990 0.6673 0.7429 0.6859
ARI 0.6687 0.6676 0.6565 0.6607 0.6170 0.6914 0.5792
ACC 0.7865 0.8373 0.7834 0.8001 0.8032 0.8160 0.7478
CV1 0.5616 0.4697 0.5441 0.5154 0.5394 0.5480 0.6385
DCV -0.0666 0.0254 -0.0491 -0.0203 -0.0444 -0.0529 -0.1434
Running time 0.2186 1.3271 0.6290 1.8517 139.5876 3.4712 0.8499
Imbalanced Iris NMI 0.0001 0.0247 0.0061 0.6414 0.6870 0.9101 0.9101
ARI -0.0053 0.0049 -0.0023 0.7421 0.7865 0.9582 0.9582
ACC 0.5250 0.5500 0.5250 0.9500 0.9583 0.9917 0.9917
CV1 0.1179 0.0471 0.0236 1.0842 1.0607 0.9664 0.9664
DCV 0.8250 0.8957 0.9192 -0.1414 -0.1179 -0.0236 -0.0236
Running time 0.0057 0.0121 0.0142 0.0257 0.1474 0.0146 0.0115
Ecoli NMI 0.6370 0.5837 0.6087 0.5999 0.6026 0.6015 0.6530
ARI 0.4953 0.4113 0.4704 0.4176 0.4468 0.6149 0.5202
ACC 0.6310 0.5863 0.6190 0.6012 0.6012 0.6964 0.6458
CV1 0.6871 0.4671 0.7517 0.4482 0.7770 1.0893 0.7001
DCV 0.4733 0.6933 0.4087 0.7123 0.3835 0.0711 0.4604
Running time 0.0255 0.2672 0.0760 0.1770 5.3315 0.1939 0.0886
Imbalanced WBDC NMI 0.0966 0.0848 0.0828 0.6846 0.3893 0.3038 0.6907
ARI 0.0337 0.0184 0.0158 0.7855 0.5156 0.3323 0.8308
ACC 0.6260 0.5827 0.5745 0.9892 0.9539 0.9051 0.9892
CV1 0.2644 0.1418 0.1188 1.3529 1.2073 1.0540 1.3069
DCV 1.0578 1.1804 1.2034 -0.0307 0.1150 0.2683 0.0153
Runing time 0.0141 0.0234 0.0094 0.0959 1.2488 0.0870 0.0535
Htru2 NMI 0.4068 0.4101 0.4076 0.0547 0.3854 0.2611 0.5869
ARI 0.6071 0.5798 0.6076 0.0217 0.5967 0.3504 0.7331
ACC 0.9366 0.9253 0.9366 0.9097 0.9428 0.8473 0.9660
CV1 1.0891 1.0125 1.0882 1.4107 1.1813 0.7805 1.2342
DCV 0.3251 0.4017 0.3260 0.0035 0.2329 0.6337 0.1800
Running time 0.0825 0.2913 0.1094 2.0558 68.3254 0.4956 0.1119

### v3

@@ VI Deep Clustering
Clustering algorithms using Euclidean distance for defining similarity face a challenge when clustering high-dimensional data, such as images, due to the distance differences between point pairs vanishes in high-dimensional space [44]. Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data to low-dimensional representation. Many deep clustering methods combine DNNs with HKM, e.g., [21, 22, 23]. However, this design is less favorable to imbalanced data [22]. In this section, we show that EKM is a better alternative for deep clustering of imbalanced data.
@@ VI-A Optimization Procedure
where $m_{k}^{n}$ is the number of samples assigned to the $k$-th cluster before the incoming data $\mathbf{x}_{n}$, controlling the learning rate of the $k$-th centroid. Overall, the optimization procedure of DCN alternates between updating network parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (47) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (48) and (49).
@@ VI-B Clustering Performance on MNIST
To implement DCN, we refer to the code one of its authors provided, available at https://github.com/boyangumn/DCN-New. We use the default neural network structure and hyperparameters. In particular, the dimension of the low-dimensional representation is set to ten, and the parameter $\beta$ is set to one. The smoothing parameter $\alpha$ is tuned on the representation obtained by the initialized DCN network according to the strategy introduced in Section IV-C.
We first evaluate the algorithm’s performance on a balanced dataset. The considered dataset is the full MNIST [45], which contains 70,000 gray images of handwritten digits from 0 to 9. Each digit has approximately 7,000 images and each image has $28\times 28=784$ pixels. We set the smoothing parameter of EKM to $\alpha=5\mathrm{e}-3$. The clustering results are presented in TABLE VI. We compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE). SAE is a specific version of DCN that only minimizes the reconstruction error. Thus, the learned representation by SAE does not have a clustering-friendly structure. The results show that DCN outperforms SAE, highlighting the importance of a clustering-friendly structure. We can also see that DCN+EKM performs similarly to DCN+HKM. We omit the results of DCN+FKM and DCN+MEFC due to save space. Their performance is not better than DCN+HKM.
Then we evaluate the algorithm performance on an imbalanced dataset. The considered dataset is derived from the full MNIST by removing the training images of digits 1 to 9. This imbalanced dataset (we call imbalanced MNIST) contains 15,923 images, of which approximately 7,000 are digit 0, while the remaining digits (1 to 9) each have about 1,000 images. We set the smoothing parameter of EKM to $\alpha=3.8\mathrm{e}-3$. The results are summarized in TABLE VII. We can see that DCN+EKM greatly outperforms other algorithms. This finding implies that the EKM-friendly structure is crucial for deep clustering of imbalanced data.
We map the ten-dimensional representation obtained by DCN to a two-dimensional space by t-SNE [46] for visualization. The representation learned on the full MNIST dataset is displayed in Fig. 8 and the representation learned on the imbalanced MNIST dataset is displayed in Fig. 9. The visualization results suggest that the deep representation obtained by DCN+EKM is more discriminative than that obtained by DCN+HKM as the former has clearer inter-cluster margins, especially for those obtained from the imbalanced MNIST dataset. This should be attributed to the centroid repulsion mechanism of EKM, which makes data farther away in representation space. We also observe from Fig. 9 that EKM successfully identifies the large class (digit 0) from other small classes without referring to the true labels, while HKM incorrectly divides the large class into four clusters to balance the cluster sizes.
TABLE VI: Evaluation on full MNIST
@@
TABLE VII: Evaluation on imbalanced MNIST
@@
Figure 8: The t-SNE visualization of DCN representations learned on the full MNIST dataset. (a) Representations obtained by DCN+HKM, colored by reference labels. (b) Representations obtained by DCN+HKM, colored by labels generated by HKM. (c) Representations obtained by DCN+EKM, colored by reference labels. (d) Representations obtained by DCN+EKM, colored by labels generated by EKM.
Figure 9: The t-SNE visualization of DCN representations learned on the imbalanced MNIST dataset. (a) Representations obtained by DCN+HKM, colored by reference labels. (b) Representations obtained by DCN+HKM, colored by labels generated by HKM. (c) Representations obtained by DCN+EKM, colored by reference labels. (d) Representations obtained by DCN+EKM, colored by labels generated by EKM.
@@ VII Conclusion
This paper presents equilibrium K-means (EKM), a novel clustering algorithm with a new centroid repulsion mechanism effective for imbalanced data. EKM is simple, interpretable, and scalable to large datasets. Experimental results on datasets from various domains show that EKM outperforms HKM, FKM, and other state-of-the-art centroid-based algorithms on imbalanced datasets and performs comparably on balanced datasets. We also demonstrate that EKM is a better alternative to HKM and FKM in deep clustering when dealing with imbalanced data. Furthermore, we reformulate HKM, FKM, and EKM in a general form of gradient descent. We encourage readers to study more properties of this general form. Combining EKM with kernels, multi-prototype mechanisms, and other techniques to deal with more complex data structures is also an important research direction. We believe that these explorations will promote the development of EKM and further solve the data imbalance issue.
@@ Appendix A Proof of Theorem 1
For concise and tidy, we denote the partial derivative $\frac{\partial h}{\partial d_{k}}(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$ as $\frac{\partial h}{\partial d_{k}}|_{n,{\tau}}$, $h(d_{1n},\cdots,d_{Kn})$ as $h$, and $h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$ as $h^{(\tau)}$. Since $h$ is a concave function at $[0,\,+\infty)^{K}$, we have
$h\leq h^{(\tau)}+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)}),$ (51)
@@
$J(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})\leq\sum_{n=1}^{N}\bigg{[}h^{(\tau)}+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)})\bigg{]}\\ =J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})+\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\bigg{(}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot\big{(}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}\\ -\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\big{)}\bigg{)}.$ (52)